Atomos Test Generator Free PYQs ExamsMy practice Log in / Sign up free

Probability (Class 12)

JEE Main 2026 (Online) 28th January Evening ShiftMCQMedium📈 Topic in 5% papers+4−1
Ques. 1JEE Main 2026, 28 January (Evening Shift)

The probability distribution of a random variable X is given below : array|c|c|c|c|c|c|c|c|c| X & 4k & 307k & 327k & 347k & 367k & 387k & 407k & 6k \ P(X) & 215 & 115 & 215 & 15 & 115 & 215 & 15 & 115 \ array If E(X) = 26315, then P(X < 20) is equal to :

  1. 35
  2. 1415
  3. 815
  4. 1115

Answer

(4) 1115

Solution

Check: 2 + 1 + 2 + 3 + 1 + 2 + 3 + 115 = 1

E(X) = xi pi = k105(56 + 30 + 64 + 102 + 36 + 76 + 120 + 42) = 526k105

526k105 = 26315

k = 263 10515 526 = 72

so the values of X are 14, 15, 16, 17, 18, 19, 20, 21, and

P(X < 20) = 1 - P(X = 20) - P(X = 21)

= 1 - 315 - 115 = 1115

Answer: (4)

Keep practising — Probability (Class 12)

Ques. 2JEE Main 2026, 24 January (Morning Shift)

From a lot containing 10 defective and 90 non-defective bulbs, 8 bulbs are selected one by one with replacement. Then the probability of getting at least 7 defective bulbs is

  1. 73108
  2. 67108
  3. 7107
  4. 81108
Ques. 3JEE Main 2026, 23 January (Evening Shift)

Bag A contains 9 white and 8 black balls, while bag B contains 6 white and 4 black balls. One ball is randomly picked up from the bag B and mixed up with the balls in the bag A. Then a ball is randomly drawn from the bag A. If the probability, that the ball drawn is white, is pq, (p, q) = 1, then p + q is equal to

  1. 24
  2. 22
  3. 23
  4. 21
Ques. 4JEE Main 2026, 22 January (Morning Shift)

If a random variable x has the probability distribution arrayccccccccc x & 0 & 1 & 2 & 3 & 4 & 5 & 6 & 7 \ P(x) & 0 & 2k & k & 3k & 2k2 & 2k & k2 + k & 7k2 array then P(3 < x 6) is equal to

  1. 0.34
  2. 0.64
  3. 0.22
  4. 0.33
Ques. 5JEE Main 2026, 21 January (Evening Shift)

A random variable X takes values 0, 1, 2, 3 with probabilities 2a+130, 8a-130, 4a+130, b respectively, where a, b R. Let and respectively be the mean and standard deviation of X such that 2 + 2 = 2. Then ab is equal to:

  1. 12
  2. 60
  3. 30
  4. 3
Ques. 6JEE Main 2025, 8 April (Evening Shift)

If A and B are two events such that P(A) = 0.7, P(B) = 0.4 and P(A B) = 0.5, where B denotes the complement of B, then P (B (A B) ) is equal to

  1. 13
  2. 12
  3. 14
  4. 16
Ques. 7JEE Main 2025, 7 April (Evening Shift)

A bag contains 19 unbiased coins and one coin with head on both sides. One coin drawn at random is tossed and head turns up. If the probability that the drawn coin was unbiased, is mn, (m, n) = 1, then n2 - m2 is equal to :

  1. 64
  2. 80
  3. 60
  4. 72
Ques. 8JEE Main 2025, 7 April (Evening Shift)

Let a random variable X take values 0, 1, 2, 3 with P(X = 0) = P(X = 1) = p, P(X = 2) = P(X = 3) and E(X2) = 2E(X). Then the value of 8p - 1 is :

  1. 2
  2. 0
  3. 3
  4. 1
Ques. 9JEE Main 2025, 4 April (Morning Shift)

A box contains 10 pens of which 3 are defective. A sample of 2 pens is drawn at random and let X denote the number of defective pens. Then the variance of X is

  1. 1115
  2. 215
  3. 35
  4. 2875
Ques. 10JEE Main 2025, 4 April (Evening Shift)

A card from a pack of 52 cards is lost. From the remaining 51 cards, n cards are drawn and are found to be spades. If the probability of the lost card to be a spade is 1150, then n is equal to .

Ques. 11JEE Main 2025, 3 April (Evening Shift)

If the probability that the random variable X takes the value x is given by P(X = x) = k(x + 1)3-x, x = 0, 1, 2, 3, , where k is a constant, then P(X 3) is equal to

  1. 19
  2. 827
  3. 727
  4. 49

All Probability (Class 12) PYQs →

PDFDownload Probability (Class 12) PYQs as PDF Free, with answer key and a QR code to every solution — print it or study offline⬇
All JEE Main Mathematics chapters